Given reaction:

C2H5OH(aq) + O2(g) —> CH3COOH(aq) + H2O(l)

The following data may be useful:

C2H5OH(aq): ΔHf° = -288.3 kJ/mol; S° = 148.5 J/mol•K

CH3COOH(aq): ΔHf° = -485.8 kJ/mol; S° = 178.7 J/mol•K

(a) Calculate ΔH° and ΔS° for this process.

(b) Is the reaction spontaneous at 25°C? At 4°C?

(c) The heat of vaporization for acetic acid is 24.3 kJ/mol. Its normal boiling point is 118.5°C. Calculate ΔS° for the reaction: CH3COOH(l) —> CH3COOH(g)

(d) What is the standard molar entropy for CH3COOH(g), taking S° for CH3COOH(l) to be 159.8 J/mol•K

(e) Calculate ΔG at 25°C for the formation of acetic acid from ethanol (see reaction above) when [CH3COOH] = 0.200 M, PO2 = 1.13 atm, and [C2H5OH] = 0.125 M.

(f) Calculate ΔG° for the ionization of acetic acid at 25°C (Ka = 1.8 x 10-5).


1
Expert's answer
2021-03-22T06:33:10-0400

"a) \u0394H\u00b0=\u0394Hf\u00b0"

H=-288.3--485.8=197.5kj/mol

ΔS°=178.7-148.5 =30.2kj/mol

b)The reaction becomes spontaneous at 250C.

c)"S= Hv\/T"

S=24.3/392 =0.062

d)S° for CH3COOH(l) to be 159.8 J/mol•K

e)CH3COOH] = 0.200 M, PO2 = 1.13 atm, and [C2H5OH] = 0.125 M.

f)"G= -RTln Ka"

=-8.314×298.15ln1.8×10^-5

=27081.44